# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ 8.5 Label deficiency
Lastly, the concept of federated transfer learning (Fig.  6C.) can be categorized into two cases: homogeneous and heterogeneous. Homogeneous transfer learning can be further classified into single- and multi-source architectures, depending on the number of source and target domains involved [Jin et al., 2023]. In a single-source architecture, there is a central server that maintains labeled data and clients with unannotated data. Since each institution has different data distribution (i.e., the data is heterogeneous across the involved sites), it creates a unique target domain, requiring adaptation methods to tackle multiple targets. The multi-source setting is much more complicated as it involves multiple source domains with labeled data transferring knowledge to a single unlabeled target domain. In this case, selecting a source domain with valuable knowledge is required without directly observing the source data. On the other hand, to make the heterogeneous transfer learning work, there must be cross-domain links between data to bridge the gap between the different feature spaces (e.g., cross-domain links between imaging modalities used for scanning the patient).
@@ 8.6 System architecture
Many institutions lack solid computing resources and reliable network infrastructure [Darzidehkalani et al., 2022]. Hence, redundant computing facilities and data centers are critical to prevent data loss. An effort in architecture planning should be made to handle the failure of one or more centers without disrupting the overall training process. The FL system must be adaptable to operate with various centers in case the institution is no longer involved in the training and maintaining the global model to solve a particular MI task.
@@ 9 Conclusions
This article thoroughly reviewed the latest developments in federated learning, expressly focusing on medical image analysis applications. The paper classified newly introduced aggregation and learning techniques specifically designed to address the challenges attributed to federated learning that differentiate it from standard deep learning pipelines. These challenges are incredibly significant in the case of multi-institutional medical imaging data delivered with different equipment and/or vendors and cover issues such as accuracy of the jointly trained model, data heterogeneity, label deficiency, the occurrence of malicious clients, data security, system architecture, and communication efficiency across the involved sites. The study has inspected open-source software packages that allow for rapid prototyping of a deep learning algorithm in a federated way using various computational supplies. We believe that the federated learning concept is a giant step in demarcating the future of deep learning involving multi-institutional medical imaging data.

### v2

@@ 9.5 Label deficiency
Lastly, the concept of federated transfer learning (Fig.  7C.) can be categorized into two cases: homogeneous and heterogeneous. Homogeneous transfer learning can be further classified into single- and multi-source architectures, depending on the number of source and target domains involved [Jin et al., 2023]. In a single-source architecture, there is a central server that maintains labeled data and clients with unannotated data. Since each institution has different data distribution (i.e., the data is heterogeneous across the involved sites), it creates a unique target domain, requiring adaptation methods to tackle multiple targets. The multi-source setting is much more complicated as it involves multiple source domains with labeled data transferring knowledge to a single unlabeled target domain. In this case, selecting a source domain with valuable knowledge is required without directly observing the source data. On the other hand, to make the heterogeneous transfer learning work, there must be cross-domain links between data to bridge the gap between the different feature spaces (e.g., cross-domain links between imaging modalities used for scanning the patient).
@@ 9.6 System architecture
Many institutions lack solid computing resources and reliable network infrastructure [Darzidehkalani et al., 2022b]. Hence, redundant computing facilities and data centers are critical to prevent data loss. An effort in architecture planning should be made to handle the failure of one or more centers without disrupting the overall training process. The FL system must be adaptable to operate with various centers in case the institution is no longer involved in the training and maintaining the global model to solve a particular MI task.
9.7 Model heterogeneity
The infrastructures involved in a FL-based system typically provide different computational resources, leading to the choice of neural network architectures tailored to specific hardware capabilities or application needs [Afonin and Karimireddy, 2022, Nezhad et al., 2025, Weng et al., 2025]. Attempting to average the parameters of models with different architectures using standard FL methods, such as FedAvg, often fails. Even when models enclose similar structures, challenges can arise due to the non-convex loss landscapes [Afonin and Karimireddy, 2022]. Consequently, there is a demand to shift the paradigm from merely transmitting raw parameters to sharing derived knowledge. The two most commonly used strategies for handling this issue are knowledge distillation [Weng et al., 2025] and generative methods [Nezhad et al., 2025].
Knowledge distillation protocols employ the model-agnostic nature of prediction outputs to transfer knowledge between models. This involves training a student model on the predictions made by a teacher model rather than aligning weights [Afonin and Karimireddy, 2022, Weng et al., 2025]. However, some methods require extra access to publicly available datasets to facilitate knowledge distillation, which may not always be accessible [Wu et al., 2022]. Further, knowledge distillation techniques suffer from sequential information loss, potentially leading to reduced model accuracy [Afonin and Karimireddy, 2022]. In contrast, methods that mitigate catastrophic forgetting often come with significantly higher computational costs [Weng et al., 2025].
An alternative is to employ generative priors collaboratively trained to capture the data distribution across multiple sites. Each client can use this prior locally to generate synthetic data. This approach enables clients to train their selected heterogeneous networks independently while benefiting from globally informed and augmented datasets [Kang et al., 2025, Nezhad et al., 2025]. However, generative methods, such as adversarial or diffusion models, often face challenges, such as unstable training, high inference or sampling costs [Nezhad et al., 2025]. Additionally, these methods may encounter scalability issues, as achieving stable convergence becomes increasingly complex with a higher number of clients [Kang et al., 2025].
To effectively address model heterogeneity, it is necessary to develop methods that are accessible to hospitals and medical centres, which often work with limited computational resources and may lack access to extensive external datasets. Data heterogeneity, discussed in section 9.2, further complicates model-agnostic FL by limiting the transferable knowledge between institutions [Afonin and Karimireddy, 2022].
@@ 10 Conclusions
This article thoroughly reviewed the latest developments in federated learning, expressly focusing on medical image analysis applications. The paper classified newly introduced aggregation and learning techniques specifically designed to address the challenges attributed to federated learning that differentiate it from standard deep learning pipelines. These challenges are incredibly significant in the case of multi-institutional medical imaging data delivered with different equipment and/or vendors and cover issues such as accuracy of the jointly trained model, data and model heterogeneities, label deficiency, the occurrence of malicious clients, data security, system architecture, and communication efficiency across the involved sites. The study has inspected open-source software packages that allow for rapid prototyping of a deep learning algorithm in a federated way using various computational supplies. Finally, we explore real-word implementations of FL-based systems and discuss the challenges in deploying these systems in clinics. We believe that the federated learning concept is a giant step in demarcating the future of deep learning involving multi-institutional medical imaging data.
Competing interests
None Declared.
Declaration of generative AI in scientific writing
The authors did not use generative AI in writing of this paper.
CRediT authorship contribution statement
Dominika Ciupek: Conceptualization, Methodology, Investigation, Resources, Software, Writing - Original Draft, Writing – review & editing, Visualization; Maciej Malawski: Conceptualization, Writing - Original Draft, Writing – review & editing, Supervision, Funding acquisition; Tomasz Pieciak: Conceptualization, Methodology, Investigation, Resources, Writing - Original Draft, Writing – review & editing, Supervision, Project administration.